Suppose that you and your friend are part of a group of 14 potential jurors. If two people from this group are to be chosen at random for the last remaining seat on the jury, what is the probability that you and your friend are the two people chosen?
Answer by stanbon(57282)

Suppose that you and your friend are part of a group of 14 potential jurors. If two people from this group are to be chosen at random for the last remaining seat on the jury, what is the probability that you and your friend are the two people chosen?
------------------------------
You and your friend constitute one pair.
---
# of possible pairs = 14C2
----
P(you and your friend are chosen) = 1/14C2 = 1/91
======================
Cheers,
Stan H.
==========